# Pair Production in time-dependent Electric field at Finite times

### v4

@@ IV.2 Longitudinal Momentum Spectrum
The time scales governing the process are influenced by the electric field strength $E_{0}$, as shown in Table II. A clear pattern emerges—these time scales consistently follow a specific trend as $E_{0}$ increases. Notably, for a fixed pulse duration $\tau$, key events occur earlier in time with increasing $E_{0}$. In general, these time scales can be quantified in terms of the electric field parameters, specifically $E_{0}$ and $\tau.$ Although a detailed analysis of time scales is beyond the scope of this paper, they can be expressed as:
$t_{cp}=0.1698\,\tau^{1.102}{E_{0}}^{-1.478},\\ t_{sep}=0.456\,\tau^{1.065}{E_{0}}^{-1.059},\\ t_{dis}=1.831\,\tau^{0.871}{E_{0}}^{-0.830}.$ (85)
These relations highlight the nonlinear dependence of each timescale on both the pulse duration and the electric field strength. Our analysis is specifically applicable to Sauter pulsed electric fields within the range $305<\tau<$ 30 and $0.50.1<E_{0}<$ 0.5. The time scales found in ref. Diez et al. (2023) differ from the scaling we have identified in our analysis.
We have not yet discussed the Keldysh parameter, $\gamma$, which classifies the pair production process as either a multi-photon or tunneling mechanism. Additionally, we are investigating whether the oscillatory behavior explicitly depends on $\gamma$. For this study, we select two different combinations of $E_{0}$ and $\tau$ to vary $\gamma$ and examine its behavior over a finite time interval. Specifically, we set $\gamma=1$, corresponding to two distinct parameter configurations $(E_{0},\tau)$ associated with the Sauter pulse electric field. It is known that $\gamma=1$ represents an intermediate regime involving both tunneling and multi-photon processes, which are inherently nonperturbative Avetissian et al. (2002). This approach allows us to demonstrate how the same $\gamma$ value can lead to qualitatively different outcomes in the time evolution of the pair production process.
Figure 6 shows LMS for the Sauter pulse electric field with $E_{0}=0.2E_{c}$ and $\tau=5\,[m^{-1}]$, while Fig.7 shows LMS for field parameter $E_{0}=0.1E_{c}$ and $\tau=10\,[m^{-1}]$. From Figs. 6 (a) to (c) we can see that due to the electric field smooth unimodal peaked profile accelerated towards negative $z-$ direction and after the pulse duration $t=6[m^{-1}]$ we observed smooth profile becomes deformed near the region $-0.5<p_{\parallel}<0.5.$ At $t=11.5[m^{-1}]$, the interference effect is observed, due to which the smooth unimodal peak structure becomes a multi-modal structure as shown in Figs.6(d) and 6(e). That interference effect nearly disappears at $t\approx 4\tau.$ Comparing Figs. 6(d), (e) and 7(d), (e), clear structural differences emerge. These differences give rise to two qualitatively distinct behaviors at finite times under the condition $\gamma=1$. Notably, the primary factor influencing this behavior at finite times is the pulse duration $tau$.
@@ IV.2.1 Approximate expression for longitudinal momentum distribution function
In this subsection, we aim to analyze the oscillatory behavior of the LMS in the late-time limit, specifically for $t>2\tau$. We utilize the previously derived approximate analytical expression for the momentum distribution function Eq.(78). By setting $p_{\perp}=0$, we derive an approximate expression that depends solely on the longitudinal momentum. We express the longitudinal momentum distribution function $f(p_{\parallel},y)$ as a series in terms of $(1-y)$ up to the second order and neglecting higher orders:
$f(p_{\parallel},y)$ $=|N^{+}(p_{\parallel})|^{2}\Bigl{(}\mathrm{C}_{0}(p_{\parallel},y)+(1-y)\mathrm{C}_{1}(p_{\parallel},y)+(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)\Bigr{)}.$ (86)
@@
In Figure 8, we illustrate the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ and the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$, derived from the longitudinal momentum distribution function given by Eq. (86). These terms are plotted as functions of longitudinal momentum at different times, allowing for a comparison of their behavior over time. Fig. 8(a), we observe that during earlier times, the term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ is more pronounced compared to $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$. The profile of $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ shows a prominent peak around $p_{\parallel}\simeq-2[m]$, largely influenced by $|\Gamma_{1}|^{2}$ (refer to Eq. (89)), which peaks near $p_{\parallel}\approx-E_{0}\tau$ depending on parameters $E_{0}$ and $\tau$. The oscillations observed within $-1<p_{\parallel}<1$ stem from the $\cos{(\Upsilon)}$ in the expression for $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$, contributing to this oscillatory pattern. Over time, the influence of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ gradually increases compared to earlier times. This term exhibits oscillations within a Gaussian envelope, with the oscillations decaying away from $p_{\parallel}\approx 0$ (see Figs. 8(b) and (c)). To understand the oscillations and their dependence on $p_{\parallel}$-value and the time variable $y$, the argument $\Upsilon$ of the cosine function is approximated as follows:
@@
$+\frac{1}{2}\tau\Bigl{(}\ln{[-16E_{0}^{2}\tau^{2}]}+\sqrt{1+4E_{0}^{2}\tau^{2}}+2E_{0}\tau\ln{[-4E_{0}\tau]}\Bigr{)},$ (91)
$K_{1}$ $=\frac{E_{0}\tau^{2}}{\sqrt{1+4E_{0}^{2}\tau^{2}}}\ln{\left[\frac{\sqrt{1+4E_{0}^{2}\tau^{2}}-1}{\sqrt{1+4E_{0}^{2}\tau^{2}}+1}\right]},$ (92)
$K_{2}$ $=\frac{\tau}{4}\Bigl{(}2-\frac{2}{1+\tau^{2}}-\frac{2}{1+4E_{0}^{2}\tau^{2}}+2\ln{[E_{0}\tau]}+2\ln{(1-y)}+\frac{1}{(1+4E_{0}^{2}\tau^{2})^{3/2}}\ln{\Bigl{[}\frac{\sqrt{1+4E_{0}^{2}\tau^{2}}-1)}{\sqrt{1+4E_{0}^{2}\tau^{2}}+1}\Bigr{]}}\Bigr{)},$ (93)
@@
According to Eq. (90), the oscillation frequency depends on $p_{\parallel}$, time variable $(1-y)$, and the field parameters $(E_{0},\tau)$. Near $p_{\parallel}=-E_{0}\tau$, the cosine oscillation flattens out, and its behavior near this point changes over time. However, regular oscillations occur at other points (see Fig. 9).
In addition to the cosine function within $\mathrm{C}_{1}(p_{\parallel},y)$, the coefficient function $\mathrm{A}_{C1}(p_{\parallel})$ also significantly influences the observed oscillatory effect
$\mathrm{A}_{C1}(p_{\parallel})$ $=\frac{4|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau}{\sqrt{1+(p_{\parallel}+E_{0}\tau)^{2}}}$ (94)
The combined effect of $\cos(\Upsilon)$ and $\mathrm{A}_{C1}(p_{\parallel})$ is observed in the spectra, we compare their behaviors by plotting them for the late-time limit, as shown in Fig.9. The function $\mathrm{A}_{C1}(p_{\parallel})$ can be approximated as a Gaussian-like profile:
$\mathrm{A}_{C1}(p_{\parallel})$ $\approx\frac{2E_{0}\tau e^{\frac{\pi\tau}{2}(2E_{0}\tau-\sqrt{1+(p_{\parallel}-E_{0}\tau)^{2}}-\sqrt{1+(p_{\parallel}+E_{0}\tau)^{2}})}}{(1+(p_{\parallel}+E_{0}\tau)^{2})\sqrt{1+\tau^{2}(1+(p_{\parallel}+E_{0}\tau)^{2})}}\Biggl{(}1+(p_{\parallel}-E_{0}\tau)^{2}$
$+\sqrt{1+(p_{\parallel}-E_{0}\tau)^{2}}(-p_{\parallel}+E_{0}\tau)\Biggr{)}$ (95)
$\frac{\partial{\mathrm{A}_{C1}}}{\partial{p_{\parallel}}}\Bigg{|}_{E_{0}=0.2,\tau=10}=0$ (96)
This leads us to find that the maxima occur at approximately $p_{\parallel}\approx-0.6[m]$. The function exhibits a nearly Gaussian-like structure.
@@
Figure 9 illustrates that the function $\mathrm{C}_{1}(p_{\parallel},y)$ combines a smooth Gaussian-like envelope, $\mathrm{A}_{C1}(p_{\parallel})$, with an oscillatory function, $\cos(\Upsilon)$. This function displays a clear pattern where the amplitude decreases smoothly as one moves away from the peak of the Gaussian envelope. The Gaussian envelope, $\mathrm{A}_{C1}(p_{\parallel})$, ensures a gradual decline in the function’s amplitude, while the oscillatory component introduces periodic variations. The frequency and amplitude of these oscillations depend on field parameters $(E_{0},\tau)$, and $p_{\parallel}$.
At $t=4\tau$, the magnitude of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ dominates over the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ as seen in Fig.8(c). The second-order term diminishes at $t=5\tau$ (see Fig.8(d)). It’s worth noting that both orders have a $(1-y)$ factor, which contributes to the distribution function. However, this contribution gradually decreases as time progresses, becoming less important in the late-time limit $y\simeq 1$. At $t\gg 2\tau$, the onset of oscillations in the Gaussian-like structure can be explained by the zeroth and first-order terms of the power series expansion of $f(p_{\parallel},y)$.
@@
On careful examination of the individual components of the longitudinal momentum distribution function (86), we observe that during electron-positron formation at $t>2\tau$, the second-order term $\bigl{(}(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)\bigr{)}$ is responsible for the primary peak located at $p_{\parallel}\approx-2[m]$ (see Fig.5). As time progresses, this primary peak diminishes, and a secondary peak at $p_{\parallel}=0$ starts to build up due to the first-order term ($(1-y)\mathrm{C}_{1}(p_{\parallel},y)$) and zeroth-order term $\mathrm{C}_{0}(p_{\parallel},y)$, which are responsible for the onset oscillation on that peak. As time progresses towards infinity, $\mathrm{C}_{1}(p_{\parallel},y)$ leads to suppression. Consequently, we observe only a secondary peak due to the dominance of the $\mathrm{C}_{0}(p_{\parallel},y)$ zeroth-order term( see Fig.5). Since the component $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ represents an oscillatory finite function whose magnitude depends on $t$, the magnitude of this function plays a crucial role in determining the dynamics of $f(p_{\parallel},t)$ in $p_{\parallel}$-space at finite times.
Now, we find out an asymptotic expression for the distribution function in the limit $y\approx 1$. At this stage, the second and first-order terms become negligible, leaving only the zeroth-order term $\mathrm{C}_{0}(p_{\parallel},y\rightarrow 1)$ surviving:
$f(p_{\parallel},y\rightarrow 1)$ $=|N^{+}(p_{\parallel})|^{2}\mathrm{C}_{0}(p_{\parallel},y\rightarrow 1)$
$=4\omega_{1}|N^{+}(p_{\parallel})|^{2}(\omega_{1}+P_{1})|\Gamma_{1}|^{2}$ (97)
@@
$f(p_{\parallel})$ $\approx 2e^{2\pi\tau\bigl{(}E_{0}\tau-\sqrt{1+E_{0}^{2}\tau^{2}}\bigr{)}}\exp{\Bigl{\{}-\pi\frac{p_{\parallel}^{2}\tau}{(1+E_{0}^{2}\tau^{2})^{3/2}}\Bigr{\}}}.$ (99)
@@ IV.2.2 LMS in the multi-photon regime
An interesting qualitative contrast becomes evident when comparing the previous scenario with $\gamma=0.5$ to the current situation. The presented figure illustrates this distinct behavior. Notably, a multi-modal pattern emerges at a specific moment, characterized by the presence of more than two peaks, occurring approximately at $3\tau$.
@@
$\mathrm{C}_{1}(p_{\parallel},y)$ $\approx\mathrm{A}_{C1}(p_{\parallel})\cos{(\Upsilon)},$ (101)
with,
$\mathrm{A}_{C1}(p_{\parallel})$ $\approx\frac{1}{2}e^{\frac{\pi\tau}{2}(E_{0}\tau(2-E_{0}\tau)-2)}e^{p_{\parallel}^{2}(1-E_{0}^{2}\tau^{2})}(4+2E_{0}\tau(2-E_{0}\tau-2E_{0}^{2}\tau^{2}))+p_{\parallel}^{2}(-2+E_{0}\tau(12+57E_{0}\tau))$
$+2p_{\parallel}(-2-10E_{0}\tau-6E_{0}^{2}\tau^{2}+15E_{0}^{3}\tau^{3})$ (102)
@@
$+p_{\parallel}\frac{E_{0}\tau^{2}}{\sqrt{1+E_{0}^{2}\tau^{2}}}\ln{\Biggl{(}\frac{E_{0}^{2}\tau^{2}(1-y)}{(1+E_{0}^{2}\tau^{2})}\Biggr{)}}+\frac{p_{\parallel}^{2}\tau}{2(1+E_{0}^{2}\tau^{2})^{3/2}}\Biggl{(}-2E_{0}^{2}\tau^{2}+\ln{\Biggl{(}\frac{(1-y)}{(1+E_{0}^{2}\tau^{2})}\Biggr{)}}\Biggr{)}$ (103)
In Fig. 12, for the given field parameters, a decrease in the oscillation frequency of the $\cos{(\Upsilon)}$ function is observed compared to the previous case with $\gamma=0.5$. As the oscillation frequency decreases for smaller values of $\tau$, this is reflected in the momentum Spectrum. Near $p_{\parallel}=-E_{0}\tau$, the cosine function behaves irregularly, flattening out and oscillating regularly away from this point. The function $\mathrm{A}_{C1}(p_{\parallel})$ plays a crucial role here, resembling a Gaussian single-peak structure with its peak located at $p_{\parallel}=-0.1[m]$, as represented by the brown curve in Fig. 12. The width of the Gaussian envelope and its peak location affect the overall behavior of $\mathrm{C}_{1}(p_{\parallel},y)$. Since the field parameters $(E_{0},\tau)$ influence the occurrence of the peak in the function $\mathrm{A}_{C1}$, they also affect the behavior of the $\cos{(\Upsilon)}$ function. Consequently, the presence of oscillations within the envelope depends on the field parameters. The appearance of multiple peaks in the spectrum is the combined effect of the oscillation of the cosine function and the function $\mathrm{A}_{C1}(p_{\parallel})$. As seen in Fig. 12, the black dotted curve represents $C_{1}(p_{\parallel},y)$, which shows fewer irregular oscillations. In the long-time limit, the first-order term contribution becomes suppressed, and only the zeroth-order term, $C_{0}(p_{\parallel},y)$, becomes dominant and governs $f(p_{\parallel},y\rightarrow 1).$ As a result, the multi-modal pattern gradually dissipates, eventually leading to a singular smooth peak above $t>5\tau$.

### v5

@@ IV.2 Longitudinal Momentum Spectrum
The time scales governing the process are influenced by the electric field strength $E_{0}$, as shown in Table II. A clear pattern emerges—these time scales consistently follow a specific trend as $E_{0}$ increases. Notably, for a fixed pulse duration $\tau$, key events occur earlier in time with increasing $E_{0}$.
We have not yet discussed the Keldysh parameter, $\gamma$, which classifies the pair production process as either a multi-photon or tunneling mechanism. Additionally, we are investigating whether the oscillatory behavior explicitly depends on $\gamma$. For this study, we select two different combinations of $E_{0}$ and $\tau$ to vary $\gamma$ and examine its behavior over a finite time interval. Specifically, we set $\gamma=1$, corresponding to two distinct parameter configurations $(E_{0},\tau)$ associated with the Sauter pulse electric field. It is known that $\gamma=1$ represents an intermediate regime involving both tunneling and multi-photon processes, which are inherently non-perturbative Avetissian et al. (2002). This approach allows us to demonstrate how the same $\gamma$ value can lead to qualitatively different outcomes in the time evolution of the pair production process. Figure 6 shows LMS for the Sauter pulse electric field with $E_{0}=0.2E_{c}$ and $\tau=5\,[m^{-1}]$, while Fig.7 shows LMS for field parameter $E_{0}=0.1E_{c}$ and $\tau=10\,[m^{-1}]$. From Figs. 6 (a) to (c) we can see that due to the electric field smooth unimodal peaked profile accelerated towards negative $z-$ direction and after the pulse duration $t=6[m^{-1}]$ we observed smooth profile becomes deformed near the region $-0.5<p_{\parallel}<0.5.$ At $t=11.5[m^{-1}]$, the interference effect is observed, due to which the smooth unimodal peak structure becomes a multi-modal structure as shown in Figs.6(d) and 6(e). That interference effect nearly disappears at $t\approx 4\tau.$
Comparing Figs. 6(d), (e) and 7(d), (e), clear structural differences emerge. These differences give rise to two qualitatively distinct behaviors at finite times under the condition $\gamma=1$. Notably, the primary factor influencing this behavior at finite times is the pulse duration $\tau$.
@@ IV.2.1 Approximate expression for longitudinal momentum distribution function
In this subsection, we aim to analyze the oscillatory behavior of the LMS in the late-time limit, specifically for $t>2\tau$. We utilize the previously derived approximate analytical expression for the distribution function Eq.(98). By setting $p_{\perp}=0$, we derive an approximate expression that depends solely on the longitudinal momentum and the transformed time variable $y$. The longitudinal momentum distribution function $f(p_{\parallel},y)$ is then expressed as a series expansion in powers of $(1-y)$ up to the second order and neglecting higher-order terms:
$f(p_{\parallel},y)$ $=|N^{(+)}(p_{\parallel})|^{2}\Bigl{(}\mathrm{C}_{0}(p_{\parallel},y)+(1-y)\mathrm{C}_{1}(p_{\parallel},y)+(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)\Bigr{)}.$ (109)
@@
In Figure 8, we illustrate the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ and the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$. These terms are plotted as functions of longitudinal momentum at different times, allowing for a comparison of their behavior over time. Fig. 8(a), we observe that during earlier times, the term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ is more pronounced compared to $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$. The profile of $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ shows a prominent peak around $p_{\parallel}\simeq-2[m]$, largely influenced by $|\Gamma_{1}|^{2}$ (refer to Eq. (112)), which peaks near $p_{\parallel}\approx-E_{0}\tau$ depending on parameters $E_{0}$ and $\tau$. The oscillations observed within $-1<p_{\parallel}<1$ stem from the $\cos{(\Upsilon)}$ in the expression for $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$, contributing to this oscillatory pattern. Over time, the influence of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ gradually increases compared to earlier times. This term exhibits oscillations within a Gaussian envelope, with the oscillations decaying away from $p_{\parallel}\approx 0$ (see Figs. 8(b) and (c)). To understand the oscillations and their dependence on $p_{\parallel}$-value and the time variable $y$, the argument $\Upsilon$ of the cosine function is approximated as follows:
@@
$+\frac{1}{2}\tau\left(4\ln{(2)}(1+2E_{0}\tau)+2\ln{(E_{0}\tau)}(1+E_{0}\tau)+\sqrt{1+4E_{0}^{2}\tau^{2}}\right),$ (114)
$K_{1}$ $=\frac{E_{0}\tau^{2}}{\sqrt{1+4E_{0}^{2}\tau^{2}}}\ln\left(\frac{\sqrt{1+4E_{0}^{2}\tau^{2}}-1}{\sqrt{1+4E_{0}^{2}\tau^{2}}+1}\right),$ (115)
$K_{2}$ $=\frac{\tau}{4}\left[2-\frac{2}{1+\tau^{2}}-\frac{2}{1+4E_{0}^{2}\tau^{2}}+2\ln(E_{0}\tau)+2\ln(1-y)\right.$
$\quad\left.+\frac{1}{(1+4E_{0}^{2}\tau^{2})^{3/2}}\ln\left(\frac{\sqrt{1+4E_{0}^{2}\tau^{2}}-1}{\sqrt{1+4E_{0}^{2}\tau^{2}}+1}\right)\right],$ (116)
@@
According to Eq. (113), the time-dependent oscillation frequency depends on $p_{\parallel}$,and the field parameters $(E_{0},\tau)$. Near $p_{\parallel}=-E_{0}\tau$, the cosine oscillation flattens out, and its behavior near this point changes over time. However, regular oscillations occur at other points (see Fig. 9).
In addition to the cosine function within $\mathrm{C}_{1}(p_{\parallel},y)$, the coefficient function $\mathrm{A}_{C1}$ also plays a significant role in influencing the observed oscillatory behavior in $\mathrm{C}_{1}(p_{\parallel},y)$.
$\mathrm{A}_{C1}$ $=\frac{4|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau}{\sqrt{1+(p_{\parallel}+E_{0}\tau)^{2}}}.$ (117)
The combined effect of $\cos(\Upsilon)$ and $\mathrm{A}_{C1}$ is observed in the spectra, we compare their behaviors by plotting them for the late-time limit, as shown in Fig.9. The function $\mathrm{A}_{C1}$ can be approximated as a Gaussian-like profile:
$\mathrm{A}_{C1}$ $\approx\frac{2E_{0}\tau e^{\frac{\pi\tau}{2}(2E_{0}\tau-\sqrt{1+(p_{\parallel}-E_{0}\tau)^{2}}-\sqrt{1+(p_{\parallel}+E_{0}\tau)^{2}})}}{(1+(p_{\parallel}+E_{0}\tau)^{2})\sqrt{1+\tau^{2}(1+(p_{\parallel}+E_{0}\tau)^{2})}}\Biggl{(}1+(p_{\parallel}-E_{0}\tau)^{2}$
$+\sqrt{1+(p_{\parallel}-E_{0}\tau)^{2}}(-p_{\parallel}+E_{0}\tau)\Biggr{)},$ (118)
$\frac{\partial{\mathrm{A}_{C1}}}{\partial{p_{\parallel}}}\Bigg{|}_{E_{0}=0.2,\tau=10}=0.$ (119)
This leads us to find that the maxima occur at approximately $p_{\parallel}\approx-0.6$. The function exhibits a nearly Gaussian-like structure.
@@
Figure 9 illustrates that the function $\mathrm{C}_{1}(p_{\parallel},y)$ combines a smooth Gaussian-like envelope, $\mathrm{A}_{C1}$, with an oscillatory function, $\cos(\Upsilon)$. This function displays a pattern where the amplitude decreases smoothly as one moves away from the peak of the Gaussian envelope. The Gaussian envelope ensures a gradual decline in the amplitude of function $\mathrm{C}_{1}(p_{\parallel},y)$, while the oscillatory component introduces periodic variations.
At $t=4\tau$, the magnitude of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ dominates over the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ as seen in Fig.8(c), and the second-order term diminishes at $t=5\tau$ (see Fig.8(d)). It’s worth noting that both orders have a $(1-y)$ factor, which contributes to the distribution function. However, this contribution gradually decreases as time progresses, becoming less important in the late-time limit $y\simeq 1$. At $t\gg 2\tau$, the onset of oscillations in the Gaussian-like structure can be explained by the zeroth and first-order terms of the power series expansion of $f(p_{\parallel},y)$.
@@
On careful examination of the individual components of the longitudinal momentum distribution function in Eq.(109), we observe that during electron-positron formation at $t>2\tau$, the second-order term $\bigl{(}(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)\bigr{)}$ is responsible for the primary peak located at $p_{\parallel}\approx-2$ (see Fig.5). As time progresses, this primary peak diminishes, and a secondary peak at $p_{\parallel}=0$ starts to build up due to the first-order term ($(1-y)\mathrm{C}_{1}(p_{\parallel},y)$) and zeroth-order term $\mathrm{C}_{0}(p_{\parallel},y)$, which are responsible for the onset oscillation on that peak. As time progresses towards infinity, $\mathrm{C}_{1}(p_{\parallel},y)$ leads to suppression. Consequently, we observe only a secondary peak due to the dominance of the $\mathrm{C}_{0}(p_{\parallel},y)$ zeroth-order term( see Fig.5). Since the component $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ represents an oscillatory finite function whose magnitude depends on time, the magnitude of this function plays a crucial role in determining the dynamics of $f(p_{\parallel},t)$ in $p_{\parallel}$-space at finite times.
In the limit $y\rightarrow 1$ ( or $t\rightarrow\infty$), we find an asymptotic expression for the distribution function. At this stage, the second and first-order terms become negligible, leaving only the zeroth-order term $\mathrm{C}_{0}(p_{\parallel},y\rightarrow 1)$ surviving:
$f(p_{\parallel},y\rightarrow 1)$ $=|N^{(+)}(p_{\parallel})|^{2}\mathrm{C}_{0}(p_{\parallel},y\rightarrow 1),$
$=4\omega_{1}|N^{(+)}(p_{\parallel})|^{2}(\omega_{1}+P_{1})|\Gamma_{1}|^{2},$ (120)
@@
$f(p_{\parallel})$ $\approx 2\exp{\left\{2\pi\tau\left(E_{0}\tau-\sqrt{1+E_{0}^{2}\tau^{2}}\right)-\pi\frac{p_{\parallel}^{2}\tau}{\left(1+E_{0}^{2}\tau^{2}\right)^{3/2}}\right\}}.$ (122)
@@
We next examine how these interference phenomena manifest in the multiphoton regime $(\gamma>>1)$, where distinct spectral features emerge.
@@ IV.2.2 LMS in the multi-photon regime
An interesting qualitative contrast becomes evident when comparing the previous scenario with $\gamma=0.5$ to the current situation. The presented Figs.11 illustrates this distinct behavior. Notably, a multi-modal pattern emerges at a specific moment, characterized by the presence of more than two peaks, occurring approximately at $t=3\tau$.
@@
$\mathrm{C}_{1}(p_{\parallel},y)$ $\approx\mathrm{A}_{C1}\cos{(\Upsilon)},$ (124)
with
$\mathrm{A}_{C1}$ $\approx\frac{1}{2}\exp\left\{\frac{\pi\tau}{2}\left(E_{0}\tau(2-E_{0}\tau)-2\right)+p_{\parallel}^{2}(1-E_{0}^{2}\tau^{2})\right\}[4+2E_{0}\tau(2-E_{0}\tau-2E_{0}^{2}\tau^{2})]$
$+p_{\parallel}^{2}\left(-2+E_{0}\tau(12+57E_{0}\tau)\right)+2p_{\parallel}\left(-2-10E_{0}\tau-6E_{0}^{2}\tau^{2}+15E_{0}^{3}\tau^{3}\right),$ (125)
@@
$+p_{\parallel}\frac{E_{0}\tau^{2}}{\sqrt{1+E_{0}^{2}\tau^{2}}}\ln{\Biggl{(}\frac{E_{0}^{2}\tau^{2}(1-y)}{(1+E_{0}^{2}\tau^{2})}\Biggr{)}}+\frac{p_{\parallel}^{2}\tau}{2(1+E_{0}^{2}\tau^{2})^{3/2}}\Biggl{(}-2E_{0}^{2}\tau^{2}+\ln{\Biggl{(}\frac{(1-y)}{(1+E_{0}^{2}\tau^{2})}\Biggr{)}}\Biggr{)}.$ (126)
In Fig. 12, for the given field parameters, a decrease in the oscillation frequency of the $\cos{(\Upsilon)}$ function is observed compared to the previous case with $\gamma=0.5$. As the oscillation frequency decreases for smaller values of $\tau$, this is reflected in the Spectrum. Near $p_{\parallel}=-E_{0}\tau$, the cosine function behaves irregularly, flattening out and oscillating regularly away from this point. The function $\mathrm{A}_{C1}$ plays a crucial role here, resembling a Gaussian single-peak structure with its peak located at $p_{\parallel}=-0.1[m]$, as represented by the brown curve in Fig. 12. The width of the Gaussian envelope and its peak location affect the overall behavior of $\mathrm{C}_{1}(p_{\parallel},y)$. Since the field parameters $(E_{0},\tau)$ influence the occurrence of the peak in the function $\mathrm{A}_{C1}$, they also affect the behavior of the $\cos{(\Upsilon)}$ function. Consequently, the presence of oscillations within the envelope depends on the field parameters. The appearance of multiple peaks in the spectrum is the combined effect of the oscillation of the cosine function and the function $\mathrm{A}_{C1}$. As seen in Fig. 12, the black dotted curve represents $C_{1}(p_{\parallel},y)$, which shows fewer irregular oscillations. In the long-time limit, the first-order term contribution becomes suppressed, and only the zeroth-order term, $C_{0}(p_{\parallel},y)$, becomes dominant and governs $f(p_{\parallel},y\rightarrow 1).$ As a result, the multi-modal pattern gradually dissipates, eventually leading to a singular smooth peak above $t>5\tau$.